• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Chiranjib

    Log entry time 08:06:31 on November 25, 2008

    Entry number 250649

    keyword=OWL SHIFT SUMMARY



    Owl shift "this and that"

    Shift Leader:::::::Chiranjib Dutta
    Target Operator:::::::Peter C. Bradshaw
    Third:::::::Raffaele De. Leo

    Hours we careRun IDEvents collectedCharge gatheredPolarization of the target(-/+)Comments
    00:083878 ends1M0.03657C64/67good one/its raining outside
    00:08HAPPEX 30459 ------64/67Happex DAQ restarted.Next one has to be started @04:00
    00:09--01:053879 1M0.0368C64/67good one/still raining
    01:06--02:063880 1M0.03701C64/67good one/still raining
    02:06--03:053881 1M0.03689C64/67good one/still raining
    03:05--4:013882 1M0.03686C65/66.9good one/still raining
    04:01HAPPEX 30459 ended and 30460 started ------64/67Happex DAQ restarted.Next one has to be started @08:00
    04:01--05:033883 1M0.0369C65/66.9good one/Rain stopped outside
    05:04--6:073884 1M0.0369C65/66.9good one/Beam tripped a bit more than the other runs.
    06:07--7:073885 1M0.03689C63.8/66.8good one.
    07:16--7:483886 0.5M---63.8/66.8This is a test for BigBite wire chamber High voltage.( By Xin Qian)
    7:49--8:003887 0.1M0.004463.8/66.8MCC said they wanted to do some tests which would take 5 to 10 min.So after talking to Eugene, we decided to take data until the beam stops i.e. till around 8:30A once it comes back.MCC will inform once they are done.



    COMMENTS

    This has been a pretty calm and peaceful shift.We have had 7 full 1M event runs with a total charge accumulation of ~ 0.257C. A successful crossing of 400 spin flip counts during the shift.The target behaved as normally as it usually does.The detectors really worked hard during the whole night to refrain themselves from triggering any alarm.Now its time to sleep for us !!!